# In-Context Learning Enhanced Credibility Transformer

### v1

@@ 3 Bühlmann credibility and interpretations
We establish a theoretical connection between the ICL mechanism and linear Bühlmann, (1967) credibility. In the ICL setting, we partition our data into context and target batches, ${\cal B}={\cal B}_{\rm context}\cup{\cal B}_{\rm target}$. The ICL objective is to enhance predictions on the target batch augmenting the available information by the context batch. There are multiple credibility mechanisms involved in this architecture. Besides the base Credibility Transformer, there is the obvious one giving the decorated tokens (2.4). The weights $v_{i}/(v_{i}+\kappa)$ in these decorated tokens precisely reflect linear Bühlmann, (1967) credibility that assigns a higher credibility weight for bigger case weights $v_{i}$. E.g., in claims counts modeling this is used to balance the different exposure lengths between the instances for experience rating. The following proposition unveils a more hidden credibility.
@@
This explanation (3.1) looks rather convincing, however, it hides an essential feature that makes this consideration different from linear Bühlmann, (1967) credibility. Namely, the responses $(Y_{j})_{j\in{\cal I}_{\rm context}}$ also enter the attention weights $(a_{i,j})_{j\in{\cal I}_{\rm context}\cup\{i\}}$; not being expressed in this notation. That is, in fact, we have a non-linear structure (3.1) in the responses $(Y_{j})_{j\in{\cal I}_{\rm context}}$ and a linear Bühlmann, (1967) credibility interpretation only reflects a first order statement. This point could be modified to a proper linear credibility formula by considering the changed query and key tensors
@@
i.e., changing the queries and keys to purely feature driven (excluding the responses), would give a proper linear credibility formula in the observations $Y_{j}$ in (3.1). This option can easily be used if we have a single ICL Transformer layer, but cannot easily be adapted to deep ICL Transformers.
@@
with query tensor $\mathbf{Q}=[\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{n}]^{\top}$ and key tensor $\mathbf{K}=[\boldsymbol{k}_{1},\ldots,\boldsymbol{k}_{n}]^{\top}$. The softmax operator normalizes this to 1, thus, $(a_{i,j})_{j=1}^{n}$ reflect normalized weights. If we select an instance from the target set $i\in{\cal I}$, we receive through the masking operator $\mathbf{M}$
@@ 4 Numerical example
We evaluate our approach on the well-known French MTPL dataset of Dutang et al., (2024), which has been widely studies in the recent actuarial literature. The dataset contains policy-level information with claim frequencies, allowing us to assess the proposed ICL-Credibility Transformer model’s ability to predict insurance claim rates. The characteristics of the dataset are summarized in Table 1.555We use the data-cleaned version that can be downloaded from https://aitools4actuaries.com/.
@@
Exposure: Exposure (in years)
The dataset exhibits typical characteristics of insurance portfolios: highly imbalanced with most policies having zero claims, heterogeneous exposure periods, and a mix of categorical and continuous risk factors.
We use the Poisson deviance loss as our primary loss metric, consistent with actuarial practice and presenting a strictly consistent scoring function, see Gneiting and Raftery, (2007),
@@
with $\widehat{\mu}_{i}=\widehat{\mu}(\boldsymbol{x}_{i})$ being the predictor for the claim $Y_{i}$.
We report on two main versions of the ICL-Credibility Transformer. The first one is exactly the ICL-Credibility Transformer as described in the previous section with the non-linear credibility mechanism in the attention head, see (3.1). The second one is the simpler linearized ICL-Credibility Transformer, where we restrict the queries and keys to be purely feature driven (excluding the responses), see (3.2). For the former model, we utilize two ICL Transformer layers and for the latter model, we utilize one ICL Transformer layer, so as to preserve the linear credibility structure, as discussed above. The training will have three phases. In the first phase, for both of the models, we perform a single training run on the base Credibility Transformer model. For this, we select the AdamW optimizer (learning rate $10^{-3}$, weight decay $10^{-2}$, $\beta_{2}=0.95$), batch size 1024, Poisson deviance loss, and a 15% validation split with early stopping (patience 20). Then, we freeze the CT model’s decoder and insert the outcome token decorator and the ICL Transformer layers between the CLS token and the output of the Credibility Transformer model. In this second training phase, we train the ICL Transformer, base Credibility Transformers encoder and the initial embedding weights for 50 epochs with the AdamW optimizer (learning rate $3\cdot 10^{-4}$, weight decay $10^{-2}$, $\beta_{2}=0.95$) while maintaining causal masking so that target examples do not interact. In other words, we keep only the decoder frozen in this training step, and allow all other weights, including those of the CT, to be modified. Batches are constructed as a concatenation of a context set and a held-out chunk of examples, and we apply the loss only on the held-out chunk via sample weights. The final third phase performs a fine-tuning step of the entire architecture for 20 epochs, i.e., including all of the components of the ICL-Credibility Transformer (the base Credibility Transformer, the context token decorator, the ICL transformer and the decoder) with a small AdamW learning rate $3\cdot 10^{-5}$ and early stopping (patience 10).
@@
To form ${\cal B}_{\rm context}$ for each target chunk, we retrieve neighbors from the base Credibility Transformer model’s CLS token embedding space; we interpret similarity in the CLS token as similarity in risk behavior. By default, we use cosine similarity implemented as inner product on $\ell_{2}$-normalized embeddings. We query $K=64$ nearest neighbors per target example and then build a unique candidate pool by taking the union across the target chunk and ranking candidates by similarity. After ranking by similarity, we take the top $c=1000$ candidates [^6] In more detail, for each test chunk of size $m$, we retrieve the top-$K$ nearest training neighbors per test row in the base embedding space (cosine via inner product on $\ell_{2}$-normalized vectors). Let $\mathcal{N}(x_{r})$ denote the $K$ neighbors of test point $x_{r}$ and define the candidate pool $\mathcal{C}=\bigcup_{r=1}^{m}\mathcal{N}(x_{r})$. We assign each candidate $j\in\mathcal{C}$ a single score given by its best match over the chunk, namely $s_{j}=\max_{r}\mathrm{sim}(x_{r},j)$ for cosine similarity. We then sort the candidates by $s_{j}$ in descending order for the cosine similarity, break ties deterministically, and select the first $c=1000$ unique items as the context set.. We employ the FAISS package (Johnson et al.,, 2019) for high-throughput nearest neighbor search. For a summary, see Table 2.
@@
Base CT pretraining AdamW (LR $10^{-3}$, WD $10^{-2}$, $\beta_{2}=0.95$), 100 epochs, batch 1024
ICL fine-tuning AdamW (LR $3\!\cdot\!10^{-4}$, WD $10^{-2}$, $\beta_{2}=0.95$), 50 epochs
Joint fine-tuning AdamW (LR $3\!\cdot\!10^{-5}$), 20 epochs
@@
We compare the ICL-Credibility Transformer against several baselines taken from Wüthrich and Merz, (2023), Brauer, (2024), Richman et al., 2025a and Richman et al., 2025b . Table 3 presents the out-of-sample performance comparison across all models.
@@
CAFFT 27,133 23.715 ($\pm$ 0.047) 23.807 ($\pm$ 0.017)
Ensemble CAFFT 27,133 23.630 23.726
@@
Table 3: Number of parameters, in-sample and out-of-sample Poisson deviance losses (units are in $10^{-2}$). Benchmark models for Null model through Poisson GAM are taken from Wüthrich and Merz, (2023, Table 7.9); plain-vanilla FNN and ensemble results are from Wüthrich and Merz, (2023, Table 7.9); CAFFT results are from Brauer, (2024, Tables 2 and 4); Credibility Transformer results are from Richman et al., 2025a (, Table 2); Tree-like PIN results are from Richman et al., 2025b (, Table 2).
@@
Generally, because stochastic gradient descent fitting involves several items of randomness, e.g., through the random initializations of the algorithms, we perform 5 fitting runs for all architectures, and the shown results are averages over these 5 fitting runs; in brackets we provide the observed standard deviations. The single run of the base Credibility Transformer produces a result of $23.743$ (out-of-sample), which is roughly within 1 standard deviation of the results obtained in the original Credibility Transformer model $23.788\pm 0.040$. Adding the context token decorator and the ICL Transformer to the base Credibility Transformer model produces small out-of-sample improvements for the version with two layers $23.725$; however, the linearized version yields $23.766$ (out-of-sample), demonstrating a slight worsening in performance. Fine-tuning all parameters of the ICL-Credibility Transformer produces further small improvements, with the fine-tuned non-linear ICL-Credibility Transformer version achieving the best test performance $23.710$ compared to $23.723$ for the linearized version. This latter point means that the linearized version benefits from ICL once full fine-tuning is performed. The scores discussed thus far for the linearized and two-layer models are the average scores of 5 independent fitting runs (as mentioned at the beginning of this paragraph).
@@
The best scores are achieved after ensembling (averaging) the predictions of the 5 independent fitting runs together. The scores improve after ensembling the 5 models, with the linearized model achieving a score of $23.699$ before fine-tuning and $23.678$ after fine-tuning. The best out-of-sample ICL results are achieved by the two-layer model, with $23.679$ pre-fine-tuning and $23.676$ post-fine-tuning. Notably, the simpler linearized model performs almost as well as the larger non-linear model after fine-tuning.
We observe that the ICL models have quite dramatically improved upon the results of the original Credibility Transformer. Thus, we are now able to answer the remaining two questions with which we opened this paper. The ICL-Credibility Transformer improves performance on a basic Credibility Transformer model, which in itself achieves a good out-of-sample performance, both for the ICL-Credibility Transformer version with two layers and for the linearized version. Thus, the first opening question is answered in the affirmative. Moreover, we have shown that the ICL approach improves the performance of a much smaller model focusing only on a single dataset, with the 2 layer ICL-Credibility Transformer fine-tuned version achieving the best test performance. Thus, the third opening question is also answered in the affirmative. Remarkably, this means that the credibility mechanism implemented by the ICL Transformer approach taken in this paper can improve on supervised learning for actuarial applications!
We now analyze how the steps taken to build the ICL-Credibility Transformer affect the CLS tokens and the predictions. For simplicity, we select a single run of the ICL-Credibility Transformer training and perform a principal component analysis (PCA) of the CLS tokens within the test set of the base Credibility Transformer model; then, we transform the following vectors from the models using this same fitted PCA analysis: the CLS token of the original CT before modifying this model, the decorated CLS tokens before the ICL mechanism, the decorated CLS tokens after ICL augmentation, the CLS tokens after fine-tuning the ICL-Credibility Transformer end to end, the decorated CLS tokens after fine-tuning the ICL-Credibility Transformer and, finally, the decorated CLS tokens after fine-tuning the ICL-Credibility Transformer and performing ICL. Because we use the same PCA analysis for all of the models, we can directly compare the projections of the CLS tokens and the predictions. Figure 2 shows the results of this analysis for ten points selected as the deciles of the original CT frequency predictions in the test set.
Figure 2: PCA analysis of the CLS tokens within the test set of the base Credibility Transformer model; then, we transform the following vectors from the models using this same fitted PCA analysis: the CLS token of the original CT before modifying this model, the decorated CLS tokens before ICL, the decorated CLS tokens after ICL, the CLS tokens after fine-tuning the ICL-Credibility Transformer end to end, the decorated CLS tokens after fine-tuning the ICL-Credibility Transformer and, finally, the decorated CLS tokens after fine-tuning the ICL-Credibility Transformer and performing ICL. Left panel: before full end to end fine-tuning; right panel: after full end to end fine-tuning.
From the plots, we can see that the ten points undergo quite a progression from the original CLS tokens to the decorated CLS tokens after ICL, to the CLS tokens after performing ICL; fine-tuning the model end-to-end seems slightly to shrink the space inhabited by the tokens and also to shrink the most extreme predictions somewhat. To better understand this progression, we show in Figure 3 the total progression of the ten points in the PCA space, labeling the original CT predictions made for that point, the predictions after allowing the weights of the CT model to be modified and the final predictions after ICL in the fine-tuned model.
Figure 3: PCA analysis of the CLS tokens for ten points in the test set; we show the the total progression of the ten points in the PCA space, labeling the original CT predictions made for that point, the predictions after allowing the weights of the CT model to be modified and the final predictions after ICL in the fine-tuned model.
To provide a more qualitative analysis of the progression of the ten points in the PCA space, we search for the five nearest neighbors of each of the tokens of the ten points in the various different token spaces we have described above. We then compare each point with each of its neighbors to see how the predictions change. For the explicit numerical results of the following summarized description, we refer to the appendix.
Outcome decoration consistently tightens neighborhoods, with the closest-match distance typically reduced by 10-40%; for example, point 1 shrinks from $\approx 0.35$ (pre-base) to $\approx 0.30$ (pre-decorated), point 3 from $\approx 0.018$ to $\approx 0.011$, and point 9 from $\approx 0.32$ to $\approx 0.21$. The subsequent “final” stages generally broaden the local set slightly, admitting a few nearby but more heterogeneous neighbors (adjacent Regions, alternative brands, or the other fuel type) while preserving the some of the more important covariates (Area, Region, Driver Age, and Bonus-Malus). Across the points we analyze, Region and fuel coherence dominate: decoration strengthens these alignments (e.g., R11 regular drivers for point 9; R24 diesels for points 3, 4, and 10). Brand resolution also improves: decoration often pulls exact brand matches into the top set (e.g., B14 for point 2; B2 for points 3 and 10), after which the final stages allow plausible brand variants (B1/B3). The gains are largest in sparse slices: policies with rarer combinations – such as very young, high Bonus-Malus profiles or very old drivers paired with specific brands – seeing the greatest tightening after decoration and the most helpful, controlled broadening in the final stage (notably points 2 and 10).

### v2

@@ 3 Bühlmann credibility and interpretations
We establish a theoretical connection between the ICL mechanism and linear Bühlmann (1967) credibility. In the ICL setting, we partition our data into context and target batches, ${\cal B}={\cal B}_{\rm context}\cup{\cal B}_{\rm target}$. The ICL objective is to enhance predictions on the target batch by augmenting the available information by the context batch. There are multiple credibility mechanisms involved in this architecture. Besides the base Credibility Transformer, there is the obvious one giving the decorated tokens (2.4). The weights $v_{i}/(v_{i}+\kappa)$ in these decorated tokens precisely reflect linear Bühlmann (1967) credibility weights that assigns a higher credibility weight for bigger case weights $v_{i}$. E.g., in claims counts modeling this is used to balance the different exposure lengths between the instances for experience rating. The following proposition unveils a more hidden credibility mechanism.
@@
This explanation (3.1) looks rather convincing, however, this reasoning has a catch that makes it different from linear Bühlmann (1967) credibility. Namely, the responses $(Y_{j})_{j\in{\cal I}_{\rm context}}$ also enter the attention weights $(a_{i,j})_{j\in{\cal I}_{\rm context}\cup\{i\}}$; not being explicitly expressed in this notation. That is, in fact, we have a non-linear structure (3.1) in the responses $(Y_{j})_{j\in{\cal I}_{\rm context}}$ and a linear Bühlmann (1967) credibility interpretation only reflects a first order approximation. This point could be modified to a proper linear credibility formula by considering the changed query and key tensors (not depending on the responses)
@@
i.e., changing the queries and keys to purely feature driven (excluding the responses), would give a proper linear credibility formula in the observations $Y_{j}$ in (3.1). This option can easily be used if we have a single ICL Transformer layer, but cannot easily be adapted to deep ICL Transformers. We will also consider this option (3.2) in our example below.
@@
with query tensor $\mathbf{Q}=[\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{n}]^{\top}$ and key tensor $\mathbf{K}=[\boldsymbol{k}_{1},\ldots,\boldsymbol{k}_{n}]^{\top}$. The softmax operator normalizes this to 1, thus, $(a_{i,j})_{j=1}^{n}$ reflect normalized weights for all $i$. If we select an instance from the target set $i\in{\cal I}$, we receive through the masking operator $\mathbf{M}$
@@ 4 Numerical example
We evaluate our approach on the well-known French MTPL dataset of Dutang et al. (2024), which has been widely studied in the recent actuarial literature. The dataset contains policy-level information with claims frequencies, allowing us to assess the proposed ICL-Credibility Transformer model’s ability to predict insurance claim rates. The characteristics of the dataset are summarized in Table 1.555We use the data-cleaned version that can be downloaded from https://aitools4actuaries.com/.
@@
Exposure: Exposure (in yearly units)
The dataset exhibits typical characteristics of insurance portfolios: highly imbalanced with most policies having zero claims, heterogeneous exposure periods, and a mix of categorical and continuous covariates.
We use the Poisson deviance loss as our primary loss metric, consistent with actuarial practice and presenting a strictly consistent scoring function, see Gneiting and Raftery (2007),
@@
with $\widehat{\mu}_{i}=\widehat{\mu}(\boldsymbol{x}_{i})$ being the predictor for the claim $Y_{i}$ and where the square bracket is set equal to $\widehat{\mu}_{i}$ for $Y_{i}=0$.
We report on two main versions of the ICL-Credibility Transformer. The first one is exactly the ICL-Credibility Transformer as described in the previous section with the non-linear credibility mechanism in the attention head, see (3.1). The second one is the simpler linearized ICL-Credibility Transformer, where we restrict the queries and keys to be purely feature driven (excluding the responses), see (3.2). For the former model, we utilize two ICL Transformer layers and for the latter model, we utilize one ICL Transformer layer, so as to preserve the linear credibility structure, as discussed above. The training will have three phases. (Phase 1) In the first phase, for both of the models, we perform a single training run on the base Credibility Transformer model. For this, we select the AdamW optimizer (learning rate $10^{-3}$, weight decay $10^{-2}$, $\beta_{2}=0.95$), batch size 1024, Poisson deviance loss, and a 15% validation split with early stopping (patience 20). (Phase 2) Then, we freeze the Credibility Transformer’s decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ and insert the outcome token decorator and the ICL Transformer layers between the CLS token and the output of the Credibility Transformer model. In this second training phase, we train the outcome token decorator, the ICL Transformer, the base Credibility Transformer’s encoder and the initial embedding weights for 50 epochs with the AdamW optimizer (learning rate $3\cdot 10^{-4}$, weight decay $10^{-2}$, $\beta_{2}=0.95$) while maintaining causal masking so that target examples do not interact. Batches are constructed as a concatenation of a context set and a target set of examples, and we apply the loss function only on the target set via sample weights. (Phase 3) The final third phase performs a fine-tuning step of the entire architecture for 20 epochs, i.e., including all of the components of the ICL-Credibility Transformer (the base Credibility Transformer, the output token decorator, the ICL transformer and the decoder) with a small AdamW learning rate $3\cdot 10^{-5}$ and early stopping (patience 10).
@@
Base CT pre-training (phase 1) AdamW (LR $10^{-3}$, WD $10^{-2}$, $\beta_{2}=0.95$), 100 epochs, batch 1024
ICL fine-tuning (phase 2) AdamW (LR $3\!\cdot\!10^{-4}$, WD $10^{-2}$, $\beta_{2}=0.95$), 50 epochs
Joint fine-tuning (phase 3) AdamW (LR $3\!\cdot\!10^{-5}$), 20 epochs
@@
We compare the ICL-Credibility Transformer against several baselines taken from Wüthrich and Merz (2023), Brauer (2024), Richman et al. (2025a) and Richman et al. (2025b). Table 3 presents the out-of-sample performance comparison across all models.
@@
CAFTT 27,133 23.715 ($\pm$ 0.047) 23.807 ($\pm$ 0.017)
Ensemble CAFTT 27,133 23.630 23.726
@@
Table 3: Number of parameters, in-sample and out-of-sample Poisson deviance losses (units are in $10^{-2}$). Benchmark models for Null model through Poisson GAM are taken from Wüthrich and Merz (2023, Table 7.9); plain-vanilla FNN and ensemble results are from Wüthrich and Merz (2023, Table 7.9); CAFTT results are from Brauer (2024, Tables 2 and 4); Credibility Transformer results are from Richman et al. (2025a, Table 2); Tree-like PIN results are from Richman et al. (2025b, Table 2).
@@
Generally, because early stopped stochastic gradient descent fitting involves several items of randomness, e.g., through the random initializations of the algorithms, we perform 5 fitting runs for all architectures, and the shown results are averages over these 5 fitting runs; in brackets we provide the observed standard deviations. The single run of the base Credibility Transformer produces a result of $23.743$ (out-of-sample), which is roughly within 1 standard deviation of the results obtained in the original Credibility Transformer model $23.788\pm 0.040$. Adding the context token decorator and the ICL Transformer to the base Credibility Transformer model produces small out-of-sample improvements for the version with two layers $23.725$; however, the linearized version yields $23.766$ (out-of-sample), demonstrating a slight worsening in performance. Fine-tuning all parameters of the ICL-Credibility Transformer produces further small improvements, with the fine-tuned non-linear ICL-Credibility Transformer version achieving the best test performance $23.710$ compared to $23.723$ for the linearized version. This latter point means that the linearized version benefits from ICL once full fine-tuning is performed. The scores discussed thus far for the linearized and two-layer models are the average scores of 5 independent fitting runs (as mentioned at the beginning of this paragraph).
@@
The best scores are achieved after ensembling (averaging) the predictions of the 5 independent fitting runs together, that is, we average the predictions of the 5 fitting runs and we evaluate the performance of this average (ensembled) prediction. The scores improve after ensembling the 5 models, with the linearized model achieving a score of $23.699$ before fine-tuning and $23.678$ after fine-tuning. The best out-of-sample ICL results are achieved by the two-layer model, with $23.679$ pre-fine-tuning and $23.676$ post-fine-tuning. Notably, the simpler linearized model performs almost as well as the larger non-linear model after fine-tuning.
We observe that the ICL versions have quite significantly improved upon the results of the original (base) Credibility Transformer. This affirmatively answers the opening questions at the beginning of this paper, whether ICL can improve predictive models, and this means that the ICL Transformer approach taken in this paper can improve on supervised learning for actuarial applications!
We now analyze how applying the steps taken to build the ICL-Credibility Transformer affects the CLS tokens and the predictions. For simplicity, we select a single run of the ICL-Credibility Transformer training, and we perform a principal component analysis (PCA) of the CLS tokens taken from the base Credibility Transformer (i.e., before we have re-trained). The PCA is applied to the following tokens:
  • (1)
    CLS tokens of the base Credibility Transformer before modifying this model (training phase 1);
  • (2)
    (2a) CLS tokens, (2b) decorated CLS tokens before the ICL mechanism, (2c) decorated CLS tokens after performing ICL augmentation (these are shown after training phase 2);
  • (3)
    (3a) CLS tokens, (3b) decorated CLS tokens before the ICL mechanism, (3c) decorated CLS tokens after performing ICL augmentation (these are shown again after training pahase 3, i.e., fine-tuning the ICL-Credibility Transformer end-to-end).
Because we use the identical PCA for all three models (phases 1-3), we can directly compare the projections of the CLS tokens as these are being modified.
Figure 2: 2D PCA representation of the CLS tokens from the base Credibility Transformer (phase 1), before being used in the ICL model architecture. The background of the figure represents a color shade depending on the points in that region using a Voronoi-like tessellation.
Phase 1: Figure 2 shows the results of the base Credibility Transformer’s predictions for 3,000 randomly selected policies in the test set. To better understand the behavior we color code the points based on the predicted log-frequency (rate) of the policies. We then apply a Voronoi-like tessellation [^7] A Voronoi-like tessellation divides the plane into numerous smaller regions, each filled with a color representing its closest data point. to visualize how the base Credibility Transformer adjusts the risk zones. From Figure 2, we can see that the base Credibility Transformer places its perceived low-risk policies in the top-left quadrants and its perceived high-risk policies in the right-half of the plane.
Figure 3: 2D PCA representation of the (2a) CLS tokens from the base Credibility Transformer after phase 2, (2b) decorated CLS tokens before the ICL mechanism and (2c) decorated CLS tokens after performing ICL augmentation.
Phase 2: After transferring the base Credibility Transformer model to the ICL model and retraining it with the decoder frozen, we obtain the PCA plots shown in Figure 3. The CLS token representations, which previously appeared as in Figure 2, are now condensed or almost squashed together, as seen in Figure 3(a). This occurs because the base model no longer needs to encode all discriminative information itself; it can be viewed as the base transformer learning to function as part of a pipeline rather than standalone. A pattern then begins to emerge in Figure 3(b), after the decoration stage, which is largely driven by a linear projection of the embeddings. With more degrees of freedom, these two plots taken together can be interpreted as the base deferring structural differentiation to the projection layer. Importantly, the shape produced in Figure 3(b) differs from that in Figure 2, implying that it is not simply undoing the ”condensation” from (a) but rather preparing the representations for the ICL transformer.
Finally, in Figure 3(c), the points show slightly more refined adjustments. Test points attend to labeled context rows (training examples with outcome decorations), so attention pulls outcome information from similar training examples, nudging test point positions based on evidence from labeled neighbours. These small nudges make sense because the projection layer in (b) already creates a good structure; the ICL attention then makes targeted adjustments rather than complete restructuring, simply aligning cases with what has happened to similar cases. The quadrants of the Voronoi-like tessellation from Figure 3 display a more defined split between the low- and high-risk policies than those relative to the base Credibility Transformer from Figure 2; recall that the decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ is the same in both figures.
Figure 4: 2D PCA representation of the (3a) CLS tokens from the base Credibility Transformer after phase 3, (3b) decorated CLS tokens before the ICL mechanism and (3c) decorated CLS tokens after performing ICL augmentation.
Phase 3: After fine-tuning the entire model - unfreezing the decoder that was frozen in phase 2 - we obtain PCA plots very similar to those produced after phase 2. We observe similar patterns between Figure 3 and Figure 4, implying that the performance uplift comes primarily from adjusting the decoder to better exploit the more refined representations provided by the ICL transformer. By adapting to this new information, the decoder improves the model’s overall predictive power.
Figures 5 and 6 demonstrate that the ICL process does not make significant adjustments to all points. Rather, it seems to select the points that differ from the rest of their context and adjust those points based on the points surrounding context.
Figure 5: 2D PCA representation of the (4a) CLS tokens from the base Credibility Transformer after phase 2, (4b) decorated CLS tokens before the ICL mechanism and (4c) decorated CLS tokens after performing ICL augmentation, where the colors of the points represent the adjustment made on the log-scale to the base Credibility Transformer rate predictions by the ICL process.
Figure 6: 2D PCA representation of the (5a) CLS tokens from the base Credibility Transformer after phase 3, (5b) decorated CLS tokens before the ICL mechanism and (5c) decorated CLS tokens after performing ICL augmentation, where the colors of the points represent the adjustment made on the logscale to the base Credibility Transformer rate predictions by the ICL process.